STRUCTURE OF COLD, MIXED PARAHYDROGEN-DEUTERIUM CLUSTERS

\underline{RUSSELL SLITER} and ANDREY F. VILESOV, Department of Chemistry, University of Southern California, Los Angeles, CA 90089.

Vibrational spectra of para-H_{2} molecules in clusters consisting of para-H_{2} or para-H_{2}/D_{2} seeded in He have been obtained using coherent anti-Stokes Raman spectroscopy. The Q_{1}(0) line of para-H_{2}, obtained upon expansion of neat para-H_{2}/D_{2}, exhibits a high frequency vibrational shift of about 2.6 cm^{−1}with change of D_{2} concentration from 0 to 98%. This shift is assigned to a decrease of the para-H_{2} vibron band width in mixed clusters, similar to that previously observed in bulk H_{2}/D_{2} solid. 1% para-H_{2}/D_{2} clusters in He are liquid and show a much smaller shift of about 0.5 cm^{−1}and indicate phase separation into a predominantly D_{2} core and H_{2} shell. The onset of phase separation in para-H_{2}/D_{2} mixtures is predicted at approximately 3 K providing further evidence of super-cooled liquid hydrogen clusters previously studied.